a stone projected with a velocity u at an angle theta with the horizontal maximum height H1 when it is projected with velocity u at an angle (pie/2-theta) with the horizontal, it reaches maximum height H2. the relation between the horizontal range R of the projectile, Hand H2 is?

Grade:11

2 Answers

Vikas TU
7 years ago
The Range would be same but the maximum height would be different.
Hence,
H1 = u^2sin^2thetha/g
and
H2 = u^2sin^2(pi/2-thetha)/g => u^2cos^2thetha/g
H1/H2 = > tan^2thetha
This would be the relation b/w them.
Hema
5 years ago
      R=4√H1 H2=4√u2sin2theta/2g×u2sin2(90-theta)/2g=4u2sintheta costheta/2g=2u22sintheta costheta/2g =u2sin2theta/g